OpenStudy (anonymous):

Please help i dont understand at all!! As a salesperson at Roaring Waves Beach Supplies, Carissa receives a monthly base pay plus commission on all that she sells. If she sells $500 worth of merchandise in one month, she is paid $390. If she sells $1,000 of merchandise in one month, she is paid $480. Find Carissa's salary when she sells $2,400 worth of merchandise.

OpenStudy (anonymous):

i think u need to get a sistem of equations like this salary=montly base +comision 300=(x)+500y 480=x+1000y

OpenStudy (anonymous):

First find her comisssion rate. When she sold 500 dollars more of mechendise , she got 90 more dollars. SWo for every 500 dollars she sells, she gets 90 dollars. 500/90